### v2

# Muffin: Mitigating Unhelpfulness in Emotional Support Conversations with Multifaceted AI Feedback

Abstract
Emotional support conversation systems are designed to alleviate users’ emotional distress and assist them in overcoming their challenges. While previous studies have made progress, their models occasionally generate unhelpful responses, which are intended to be supportive but instead have counterproductive effects. Since unhelpful responses can hinder the effectiveness of emotional support, it is crucial to mitigate them within conversations. Our solution is motivated by two principal considerations: (1) multiple facets of emotional support are expected to be considered when developing emotional support conversation models, and (2) directly reducing the probability of generating unhelpful responses can effectively mitigate their occurrence. Accordingly, we introduce a novel model-agnostic framework named Mitigating unhelpfulness with multifaceted AI feedback for emotional support (Muffin). It first employs a multifaceted AI feedback module designed to assess the helpfulness model responses across various facets of emotional support. Leveraging contrastive learning, Muffin then reduces the unhelpful responses’ likelihoods. To validate the effectiveness of our proposed framework, we apply Muffin to various previous emotional support generation models, including the state-of-the-art. Experimental results demonstrate that Muffin can significantly mitigate unhelpful response generation while enhancing response fluency and relevance.
Muffin: Mitigating Unhelpfulness in Emotional Support Conversations
with Multifaceted AI Feedback
Jiashuo WANG, Chunpu XU, Chak Tou LEONG, Wenjie LI, Jing LI Hong Kong Polytechnic University {csjwang,cswjli,jing-amelia.li}@comp.polyu.edu.hk {chun-pu.xu,chak-tou.leong}@connect.polyu.hk
1 Introduction
Emotional support conversation systems (supporters) are designed to generate responses that can buffer the emotional distress experienced by users (help-seekers) and help users to work through the challenges they are confronting Liu et al. (2021). Recently, many studies have contributed to this field Deng et al. (2023); Zhao et al. (2023); Cheng et al. (2022); Tu et al. (2022). Despite some success, their models occasionally generate well-intended responses that produce a counterproductive support effect, i.e., exacerbating the negative emotional states of users or inhibiting effective problem-solving, as shown in Figure 1. In the psychology and communication theories, these failed support attempts are termed “unhelpful messages” Greene and Burleson (2003); Burleson (1985).
The frequency of unhelpful responses generated by some of the previous models is not extremely high, e.g, approximately of responses generated by MultiESC Cheng et al. (2022) on ESConv benchmark Liu et al. (2021) are identified as unhelpful under strict evaluation criteria. However, their occurrence can significantly undermine earlier supportive efforts and damage the trust between the help-seeker and the supporter Llewelyn et al. (1988). Therefore, mitigating models’ generation of unhelpful responses is critical. We aim to address this problem with the following two deliberations. D1 - Consideration of Multiple Facets: Many previous studies generate responses that primarily emphasize a single facet of emotional support, e.g., one of empathetic expression Li et al. (2024), communication skill efficiency Cheng et al. (2022); Liu et al. (2021), or response coherence Deng et al. (2023), in each of their models. However, such a singular emphasis on one facet often leads to the oversight of the others, potentially resulting in unhelpfulness Greene and Burleson (2003), as exemplified in Figure 1. D2 - Direct Minimization of Unhelpful Response Probability: Previous models are typically optimized by minimizing the negative log-likelihood of the golden responses. Moving beyond this optimization objective, we aim to specifically mitigate unhelpful responses by directly targeting and reducing the probability of their generation.
Accordingly, this paper introduces a novel, model-agnostic framework called Mitigating unhelpfulness with multifaceted AI feedback for emotional support (Muffin). For D1, we design a multifaceted AI feedback module within the Muffin framework. This module assesses whether a specific response is unhelpful from multiple facets of emotional support. Leveraging the advanced capabilities of recent large language models (LLMs), we implement this module by instruction-tuning LLaMA, avoiding inefficient and expensive human feedback collection. Then, we continue optimizing an emotional support conversation model with its previous training objective and a new one to implement D2. The additional objective is to minimize the likelihood of unhelpful responses, which is implemented by contrasting unhelpful responses, identified by the feedback module, and the other (non-unhelpful) ones. Through these two steps, we aim to mitigate the unhelpful responses generated by a given emotional support conversation model. Experimental results highlight the effectiveness of our framework, demonstrating that Muffin can enhance the helpfulness of previous emotional support conversation models, including those recognized as state-of-the-art. The main contributions of this work are:
1.
We recognize and address a crucial problem in recent emotional support conversation models, i.e., the generation of unhelpful responses, a key concern in effective emotional support.
2.
We propose Muffin, a novel model-agnostic framework designed to mitigate unhelpful response generation. It incorporates a multifaceted AI feedback module to distinguish unhelpful generated responses and mitigates responses identified as unhelpful by leveraging contrastive learning.
3.
We undertake experiments with the latest emotional support conversation models, including state-of-the-art ones, to demonstrate Muffin’s effectiveness in mitigating the models’ tendency to produce unhelpful responses.
2 Related Work
2.1 Emotional Support Conversation
In the domain of emotional support conversation generation, prior studies have achieved some success. Specifically, they have each emphasized and incorporated different facets of emotional support, such as empathetic expression, communication skills, and response coherence, into their respective models. Some of them consider a single facet. For example, MultiESC Cheng et al. (2022) only considers the communication skill efficacy of responses by planning response strategies. KEMI Deng et al. (2023) incorporates related external knowledge for response generation. Although the response coherence is enhanced, the efficacy of communication skills is ignored. Li et al. (2024) enhance the empathetic expression of LLMs through chain-of-though. Beyond these, MISC Tu et al. (2022) generates supportive responses considering both commonsense and communication skills, thereby enhancing two facets of emotional support. TranESC Zhao et al. (2023) incorporate commonsense, communication skills, and emotional elements into response generation. Most all these models are optimized by minimizing the negative log-likelihood of golden responses, instead of the unhelpful response likelihood.
2.2 Contrastive Learning for Text Generation
Contrastive learning was initially employed to learn meaningful representations by contrasting positive and negative samples in the field of natural language processing Logeswaran and Lee (2018); Gutmann and Hyvärinen (2012). Recently, it has been applied to text generation Jiashuo et al. (2023); Zheng et al. (2023); Liu et al. (2022), achieving impressive performance across various settings. During training, the model is exposed to a range of “hard” negative and positive samples through contrastive learning, enabling the model to distinguish preferred outputs from less desirable ones. Consequently, selecting positive and negative samples is crucial in this process. In this paper, responses generated by the model that are deemed unhelpful are negative samples, while all other responses are considered positive samples.
3 Preliminaries
3.1 Unhelpfulness of Emotional Support
We draw upon theories in psychology and communication Greene and Burleson (2003); Burleson (1985) and adopt the term “unhelpful” to characterize responses that consistently produce negative outcomes in emotional support conversations. Conversely, responses that yield positive outcomes, or at the very least, do not cause negative effects, are termed “helpful.” These theories also suggest that an unhelpful response often stems from a flaw in merely one specific facet of emotional support. Often, the flaw directly exacerbates the user’s negative feelings or hinders effective problem-solving. For instance, a response can be deemed unhelpful if it either neglects the individual’s feelings and needs (lacking empathy) or portrays the individual’s behavior as problematic (exhibiting a dysfunctional communication skill: confront). In our work, we use this feature to identify unhelpful responses.
3.2 Problem Definition
Our primary goal is to mitigate the generation of unhelpful responses. Rather than training a new model from scratch, we aim to refine a pre-trained emotional support conversation model with the dataset it was originally trained on. This process unfolds as follows. Let represent the model trained on the dataset , where denotes the model parameters. Each instance in is denoted as , with as the input and as the expected output. Usually, is the conversation context, but it contains additional related information in some models. Assume that there are samples with labels . These samples are the diverse beam search generation results of . As for the label , it represents feedback to indicate whether the sample is unhelpful () or not unhelpful (). Our objective is to refine the model’s parameters such that the likelihood of generating unhelpful samples is reduced relative to helpful ones. In this process, we only modify the training process, ensuring that the model’s architecture and the inference mechanism remain untouched. Moreover, our approach is model-agnostic. This implies that can be any deep learning model designed and trained for emotional support conversations.
4 Method
The overall framework of Muffin is outlined in Figure 2. It is composed of two principal components, each specifically designed for deliberations in Section 1: P1: Consideration of Multiple Facets and P2: Direct Mimimization of Unhelpful Response Probality, respectively. The multifaceted AI feedback module aims to identify whether a response from multiple facets of emotional support is unhelpful. The unhelpfulness mitigation module mitigates the likelihood of unhelpful responses by contrasting helpful and unhelpful responses.
4.1 Multifaceted AI Feedback
We distinguish whether a response is unhelpful from multiple facets. However, collecting feedback from humans is inefficient and costly. In addition, recent large language models (LLMs), such as the GPT series Ouyang et al. (2022) and LLaMA Touvron et al. (2023), demonstrate remarkable natural language understanding capabilities. Therefore, we decide to obtain feedback from AI.
Instruction-tuning
Prompt engineering provides a simple and straightforward approach to obtaining feedback from LLMs. However, our experiments suggest that it is challenging to manifest the full potential of LLMs for emotional support without investing significant effort in prompt design, which will be detailed later. As an alternative, we elicit the desired capabilities of the LLM via instruction tuning Wei et al. (2021). Specifically, we design task descriptions and instructions tailored to classification tasks related to different emotional support facets. We employ corresponding datasets for these tasks. The prompt employed is illustrated in Figure 3. Notably, the response class can indicate whether the response is unhelpful regarding this facet. During the training phase, all texts in italics enclosed within curly braces are provided. During inference, the model is expected to generate the response class, based on the other italicized inputs within the curly braces.
Multifaceted AI feedback module
The final feedback for an emotional support response is derived from an aggregation of AI feedback across multiple facets. For the given response and its context, we use the instruction-tuned LLM to provide feedback on all these facets respectively. If feedback from any of these facets suggests the response is unhelpful, the response is accordingly labeled as ; otherwise, the response is deemed helpful (non-unhelpful) and labeled as , as mentioned in Section 3.1.
4.2 Unhelpfulness Mitigation
We mitigate generating unhelpful responses by contrasting helpful and unhelpful responses generated by itself, which can be implemented by the following three steps:
Generating sample responses
We utilize to generate responses on its own training dataset using diverse beam search Vijayakumar et al. (2016). Thus, for each instance , there are sample responses .
Getting feedback
These responses can be generated because they have relatively high generation probabilities. However, some of them can be unhelpful responses. Therefore, we adopt the multifaceted AI feedback module to identify whether these responses are unhelpful. Thus, we obtain labels , where .
Contrasting sample responses
We expect that the model can sign a higher likelihood to the helpful responses than the unhelpful ones. Therefore, we contrast them using the following loss:
where is the margin hyperparameter. Moreover, is the length-normalized log-probability of the response , and it is be computed by:
where is the length penalty hyperparameter. In addition to the above loss, we also consider the negative log-likelihood loss to prevent the model’s generation from deviating too much from the ground truth. The loss can be formulated as:
The final loss is the combination of the above two losses:
where and are weight hyperparameters.
5 Experiments
5.1 Experimental Setups
Base models ()
Our proposed method, i.e., Muffin, is a model-agnostic approach designed to mitigate the unhelpfulness of an existing emotional support conversation model. To examine its effectiveness, we experiment with five recent models: BlenderBot (Vanilla) Roller et al. (2021), BlenderBot-Joint (Joint) Liu et al. (2021), MultiESC Cheng et al. (2022), TransESC Zhao et al. (2023), and KEMI Deng et al. (2023). We obtain each model’s parameters using its official implementation and the default hyperparameters.
When training Muffin, we try to use the same training hyperparameters as the base model , including batch size and random seed. However, we use a smaller learning rate, i.e., , to help the model converge more efficiently. We set the epoch number as since the training loss converged within one epoch. The margin hyperparamter is , the length penalty hyperparameter is , and the weight hyperparamters and are both . In the response generation phase, as described in Section 4.2, we set the number of sample responses to . Consequently, both the beam size and the number of beam groups are configured to be , while all other generation hyperparameters are the same as its base model.
ESConv dataset ()
The ESConv dataset Liu et al. (2021) is used to train the aforementioned base models. ESConv is a benchmark for emotional support conversations, comprising approximately 1K conversations with 13K utterances. All base models, with the exception of TransESC, follow the original division of ESConv for training, validation, and testing, using an 8:1:1 ratio. TransESC employs a random split while maintaining the same ratio. Notably, each model adopts different data preprocessing methods. We adhere to each base model’s specific data division and pre-processing.
5.2 Facets of Emotional Support
We consider three essential facets of emotional support: empathetic expression, skill efficiency, and response coherence, which the base models incorporate into their models. Here, we would like to describe the unhelpfulness of each facet and detail the corresponding classification dataset.
Empathetic expression
Empathetic expressions signify the supporter’s interest in and comprehension of the help-seeker’s perspective. Conversely, their absence can impede conversation engagement and obstruct establishing a trust-based relationship between the supporter and the help-seeker Morse et al. (1992). While empathy encompasses various aspects Paiva et al. (2017); Batson (2009); Bohart and Greenberg (1997), we adopt the comprehensive framework proposed by Sharma et al. (2020). It identifies three empathy communication mechanisms, i.e., emotional reaction, interpretations, and explorations, each assessed across three levels: no communication, weak communication, and strong communication. Our work considers responses exhibiting no empathy across all mechanisms unhelpful. For example, “sleeplessness can result in upsetness,” which inappropriately offers mere information, is considered unhelpful in empathetic expression, especially when responding to a statement like “I am upset”.
For training and testing LLMs in classifying unhelpful responses regarding empathetic expression, we also utilize the dataset compiled by Sharma et al. (2020). It consists of K context-response pairs. Each response within this dataset is assessed based on three previously mentioned empathy communication mechanisms. Responses consistently labeled as “no communication” across all these mechanisms are identified as unhelpful.
Skill efficiency
The applications of effective strategies can help supporters convey appropriate and impactful support messages Greene and Burleson (2003), deepening the understanding of the help-seeker’s state and facilitating problem solution Hill (2009). However, some dysfunctional skills lead to opposite effects Barsky and Coleman (2001); Burleson and Samter (1985). For example, while advice is typically seen as beneficial, advice without permission can be less effective than employing no specific skills at all. This study assesses skill efficiency using motivational interviewing skill codes Moyers et al. (2003), which include three general categories: MI adherent, MI non-adherent, and others. Responses classified as MI non-adherent category are deemed unhelpful.
We utilize the Motivational Interviewing (MI) dataset proposed by Welivita and Pu (2022) to classify unhelpful responses in the context of skill efficiency. This dataset comprises K context-response pairs, where each response is annotated with one of three MI codes: MI adherent, MI non-adherent, or others. Responses labeled as “MI non-adherent” are considered unhelpful.
Response coherence
While response coherence is a fundamental expectation in almost all conversational systems Huang et al. (2020), it holds particular significance in emotional support conversation systems. Incoherent responses can confuse and impede effective communication. We categorize responses as coherent or incoherent, with the latter labeled unhelpful.
We have synthesized a dataset, derived from the base model’s training set , specifically for the purpose of response coherence detection. Specifically, we randomly selected K context-response pairs from . For each pair, the original response is categorized as coherent. To introduce incoherence, we employ two methods: firstly, by selecting a response from a different conversation, and secondly, by modifying keywords or important information in the original response to create a subtly incoherent variant. These methods result in a total of approximately K context-response pairs, which are then utilized for classifying unhelpful responses in terms of their response coherence.
5.3 Multifaceted AI Feedback
Instruction-tuning settings
We instruction-tune a 7B LLaMA Touvron et al. (2023) for the multifaceted AI feedback module and use a low-rank adaptor (LoRA) Hu et al. (2021) for efficiency. Specifically, we freeze LLaMA’s weight and inject trainable rank decomposition matrices into query, key, value, and output layers. The learning rate is , and the training epoch is . We merge the datasets described in Section 5.2, formatting each context-response pair according to the instruction template presented in Figure 3. Consequently, this merged dataset encompasses a total of K instances, which are utilized for instruction-tuning. The dataset is partitioned into training, validation, and test sets following an 8:1:1 split ratio.
Module performance
The instruction-tuned LLaMA model is subsequently employed within the Multifaceted AI Feedback Module. The module’s effectiveness is evaluated using the instruction-tuning test set, with the results depicted in Figure 4. We report both accuracy and F1 scores for data instances throughout the entire testing set and across distinct facets. Additionally, we extend our evaluation to include the module’s performance when utilizing three other LLMs: GPT-3.5, GPT-4, and the original (vanilla) LLaMA model. Without fine-tuning, GPT-4 outperforms the other two models in terms of the classification effectiveness. Upon closer examination, GPT-3.5 is prone to classifying responses as non-unhelpful. The vanilla LLaMA model frequently classifies responses into a singular category across different facets, leading to the worst performance. However, after instruction-tuning, LLaMA exhibits a significant enhancement in performance, with an accuracy of and an F1 score of on the aggregated feedback, thereby exceeding the capabilities of its counterparts. This remarkable improvement provides strong justification for our decision to apply the 7B instruction-tuned LLaMA within the Multifaceted AI Feedback Module.
Model | B-1 | B-2 | B-3 | B-4 | R-L | METEOR | CIDEr | Extreme |
Vanilla | 18.23 | 7.02 | 3.49 | 1.99 | 16.09 | 7.31 | 14.95 | 50.48 |
Muffin | 19.43 | 7.58 | 3.66 | 2.02 | 16.26 | 7.72 | 13.90 | 51.00 |
Joint | 18.77 | 7.54 | 3.79 | 2.15 | 17.72 | 7.59 | 17.38 | 50.96 |
Muffin | 20.59 | 8.38 | 4.26 | 2.54 | 18.35 | 8.18 | 19.12 | 51.46 |
TransESC | 17.32 | 7.10 | 3.63 | 2.18 | 17.47 | 7.53 | 22.07 | 51.33 |
Muffin | 17.19 | 7.17 | 3.73 | 2.25 | 17.54 | 7.58 | 22.72 | 51.57 |
KEMI | 19.85 | 8.15 | 4.24 | 2.52 | 17.17 | 7.92 | 15.09 | 50.85 |
Muffin | 20.01 | 8.31 | 4.36 | 2.60 | 17.30 | 7.99 | 15.45 | 51.11 |
MultiESC | 21.79 | 9.19 | 4.98 | 3.05 | 20.92 | 8.93 | 28.84 | 52.59 |
Muffin | 21.83 | 9.28 | 5.12 | 3.21 | 21.26 | 8.92 | 31.26 | 52.83 |
5.4 Emotional Support Response Generation
Automatic evaluation
We evaluate the quality of emotional support responses by a range of automatic evaluation metrics, including BLEU Papineni et al. (2002) (B-1/2/3/4), ROUGE (R-L) Lin (2004), METEOR Banerjee and Lavie (2005), CIDEr Vedantam et al. (2015), and BOW Embedding-based matching score (Extreme) Liu et al. (2016). These metrics are good at evaluating the similarity between the generated response and the ground truth.
Table 1 showcases the performance of Muffin with different base models in all automatic evaluation metrics. In general, Muffin demonstrates significant enhancements across nearly all evaluation metrics. Moreover, it can be observed that the performance of Muffin is predominantly influenced by its base model , assessed through automatic evaluations.
Human evaluation
Following previous work Deng et al. (2023); Liu et al. (2021), we compare the base model and its corresponding Muffin model on five aspects, which are (1) Fluency: which model’s response is more fluent? (2) Identification: which model’s response is more skillful in identifying the user’s problem? (3) Comforting: which model’s response is better at comforting the user? (4) Suggestion: which model can give more helpful and informative suggestions? (5) Overall: which model’s response is generally more helpful from the aspect of the help-seeker? Specifically, for each -Muffin pair, we randomly select 100 instances for comparison. Then, we ask four unique human evaluators to vote which response is better. They can select “tie” if both responses are considered equal. We average their results as the final result.
Figure 5 summarizes the A/B test results on BlenderBot-Joint Liu et al. (2021), KEMI Deng et al. (2023), and MultiESC Cheng et al. (2022), along with their corresponding Muffin models. These three settings are selected for their significant performance in automatic evaluation. The inter-rater agreement, i.e., Fleiss’ Kappa Fleiss (1971), is , implying fair agreement. Our Muffin models are regarded as more helpful in general. Responses generated by Muffin models are slightly more fluent than those generated by base models. Compared with the corresponding base model , the Muffin model shows more powerful capability in identifying the help-seeker’s problem. Moreover, Muffin models can generate responses that have better effects to comfort the users than . Annotators also prefer responses generated by Muffin because of their more helpful and informative suggestions. These results prove that Muffin indeed mitigates the unhelpfulness of emotional support conversation models.
Model | B-1 | B-2 | B-3 | B-4 | R-L | METEOR | CIDEr | Extreme |
Joint | 18.77 | 7.54 | 3.79 | 2.15 | 17.72 | 7.59 | 17.38 | 50.96 |
Muffin | 20.59 | 8.38 | 4.26 | 2.54 | 18.35 | 8.18 | 19.12 | 51.46 |
Muffin (emp.) | 19.58 | 8.06 | 4.04 | 2.33 | 18.46 | 7.87 | 19.61 | 51.43 |
Muffin (skill) | 18.68 | 7.51 | 3.81 | 2.19 | 17.98 | 7.67 | 17.84 | 50.98 |
Muffin (cohr.) | 20.04 | 8.10 | 4.04 | 2.26 | 18.24 | 7.95 | 18.07 | 51.31 |
Ablation study
To assess the impact of different facets of feedback, we undertake an ablation study. Specifically, we employed AI feedback from each individual facet to distinguish helpful from unhelpful responses, subsequently contrasting them to mitigate unhelpfulness. This study is conducted using BlenderBot-Joint as the base model, given its relatively straightforward architecture and the notable improvements exhibited by Muffin.
The findings, as presented in Table 2, reveal that relying solely on AI feedback from a single facet for unhelpfulness mitigation results in diminished performance in automatic evaluations when compared to the comprehensive Muffin model. This underscores that considering multiple facets of emotional support when building emotional support conversation models (P1) is necessary. Another insight from Table 2 is that all three ablated models outperform the base model. It suggests that the proposed solution, directly mitigating unhelpful responses (P2), is reasonable and effective. Moreover, the results also indicate that the quality of helpful and unhelpful responses will influence the effects of unhelpfulness mitigation.
Conversation History | |
Help-seeker : I’m an alcoholic. I get that thought in my head and its like I become obsessed. |
Supporter : Why not try to put attention on some other things instead of a beer? |
Help-seeker : Hmmm like what? |
Model Generated Responses | |
Joint |
You can try to be normal. That is a great idea. (Lacking empathy) |
Muffin |
Yes, there are some things that you can do, but you have to be aware of your own mental health, that’s why it’s important to know you are doing the right thing. |
MultiESC |
It sounds like you have a lot on your plate and you want to be normal. (Inconsistent) I can understand how that would make you feel. |
Muffin |
I think you should focus on yourself instead of a beer. Do you have any hobbies that you enjoy? |
KEMI |
I think you should try to find some new things to do to relax. |
Muffin |
Have you tried meditation? It can help you relax and relax. |
Case study
To intuitively illustrate the superiority of Muffin over its base model, Table 3 presents a comparative case study, comparing responses generated by three prominent base models and their corresponding Muffin versions. From the comparison of BlenderBot-Joint and Muffin, we can observe that the BlenderBot-Joint implies that the help-seeker can be abnormal now. Such a statement ignores the help-seeker’s feelings, barely expressing empathy. For the comparison of responses generated by BlenderBot-Joint and Muffin, the former tends to state facts more directly, subtly implying that the help-seeker might be experiencing an abnormal state. Such a statement ignores the help-seeker’s feelings, barely expressing empathy. In contrast, Muffin conveys concern for the help-seeker’s well-being and attempts to solve the problem by shifting the help-seeker’s perspective, amplifying the empathetic undertone. In the case of MultiESC, Muffin crafts a response that aligns more closely with the context, addressing the inconsistency of the response generated by MultiESC. Lastly, comparing KEMI with Muffin, even though KEMI’s response does not exhibit glaring issues, the Muffin version stands out as more beneficial. This distinction arises because Muffin, in contrast to KEMI’s general advice, offers a more specific and actionable recommendation, aligning closely with the help-seeker’s request for precise advice.
6 Conclusion
In this work, we focus on mitigating unhelpful responses generated by recent emotional support conversation models. Such unhelpful responses, despite their well-intentioned nature, can inadvertently counteract prior supportive efforts. Analyzing the potential causes for unhelpful responses, we introduce a novel model-agnostic framework Muffin. In specific, it contains a multifaceted AI feedback module, which can discern helpful and unhelpful responses generated by a specific emotional support conversation model. Then Muffin contrasts helpful and unhelpful responses generated by this model, in order to reduce the likelihood of the unhelpful responses. Experimental results underscore Muffin’s efficacy, showcasing enhancements in both automatic evaluations and human ratings. This suggests that Muffin can mitigate helpfulness in emotional support conversations.
Limitations
Although Muffin’s effectiveness is apparent, opportunities for further refinement exist. Our present efforts address general cases of unhelpfulness, specifically targeting responses generally perceived as unhelpful by most help-seekers. Nonetheless, it is crucial to acknowledge that certain responses may adversely affect particular individuals under specific circumstances. Consequently, this underscores the need for personalizing the emotional support conversation system to meet individual user requirements. However, this personalization presents substantial challenges. Collecting and processing personal data raises serious privacy and security concerns. Striking a balance between effective personalization and the potential for privacy intrusion remains a delicate issue. Furthermore, reconciling the goals of personalizing individual user experiences with the need to generalize models across a broader user base poses a fundamental conflict in system design.
Ethical Considerations
In our experiments, we adopted open-sourced datasets, including ESConv Liu et al. (2021), empathetic classification dataset Sharma et al. (2020), and MI dataset Welivita and Pu (2022). All personally identifiable information was removed from these datasets. For the human ratings, we emphasized the comfort and well-being of our annotators. Moreover, our research explores the development of emotional support conversation systems. Compared with existing methods, our proposed approach represents a significant leap towards establishing a more secure emotional support conversation framework. Consequently, we confidently assert that our research is conducted in strict adherence to the ethical guidelines prescribed by the Association for Computational Linguistics (ACL).
Appendix A Multifacetd AI Feedback Module
This section provides more details about the multifaceted AI feedback module.
A.1 Multiple Facets of Emotional Support
Facet |
Context |
Unhelpful Response |
Empathetic Expression |
Help-seeker: Just wanna sleep. Can’t motivate at all. I don’t know what to do. |
Supporter: As Nas says, sleep is the cousin of death. It’s the kindest form of oblivion. |
Help-seeker: What do you want to be when you grow up? I don’t know what to do with my life or what I truly love when all I really want is not to exist. |
Supporter: But what does fully grown up mean? I have my own house and a job etc. Am I fully grown up now? |
Communication Skill Efficiency |
Help-Seeker: I do nothing but think about how I’m worthless and how I shouldn’t be here. I’ve never tried or contemplated suicide. I’ve always wanted to fix my issues, but I never get around to it. How can I change my feeling of being worthless to everyone? |
Supporter: Still, you should watch out for this because it can be a sign of worsening depression. |
Help-Seeker: I try to have a “heart-to-heart” with my mom, but it usually winds up being messy and frustrating afterward. Thank you for your advice and words of encouragement. Sorry if that sounds silly. |
Supporter: Stop being so afraid…you really have nothing to lose :) Well, you are right. |
Response Coherence |
Help-seeker: Is it afternoon where you are? I’m still waking up; I should have more coffee. And I am just always tired of everything these days. |
Supporter: Could you please explain why you are enjoying yourself? |
Supporter: Yes, what happened? Why are you tired? Please can you explain? Everything going ok? |
Help-seeker: I think it’s my depression. I have a very hard time getting motivated to do anything. Nothing seems worth it. |
Help-seeker: I am upset. |
Supporter: Everyone can be happy. Can you tell me more? |
For further clarification and better understanding, we have provided additional examples of unhelpful response instances related to the emotional support facets we emphasize in our work. These examples can be found in Table 4. Furthermore, we have expanded on the explanations of helpfulness and unhelpfulness in relation to each facet and the corresponding classification dataset.
Empathetic Expression
The empathetic expression framework introduced by Sharma et al. (2020) encompasses three empathy communication mechanisms, each illustrating a different approach to conveying empathy:
Emotional Reaction: This mechanism involves expressing warmth, compassion, and concern for the help-seeker. An example is the statement “I am here to help you,” which demonstrates empathy through emotional reaction.
Interpretations: This approach reflects an understanding of feelings and experiences as inferred from the help-seeker’s statements. For instance, “I understand that you feel sad” exemplifies the supporter’s interpretations.
Explorations: Responses in this category explore feelings and experiences not explicitly stated by the help-seeker. An example is the close question “Why are you feeling alone right now?”
In the framework and associated dataset outlined by Sharma et al. (2020), responses that solely offer advice (e.g., “ask your school counselor what resources they can provide”), merely present factual information (e.g., “mindful meditation helps overcome anxiety”), or are offensive or abusive (e.g., “I don’t know what to do; I am also feeling suicidal right now”) are considered non-empathetic and characterized as “no communication.” However, as noted by Greene and Burleson (2003), information and suggestions can often be helpful, particularly when the help-seeker requires. Consequently, in our work, responses containing pertinent information or suggestions are classified as “weak/strong communication” if they align with the help-seeker’s request in the context. To identify such responses, we utilize GPT-4 to detect contexts where the help-seeker explicitly (e.g., “any suggestions”) or implicitly (e.g., “I don’t know what to do”) requests information or suggestions. Subsequently, these responses are manually evaluated to determine if they provide reasonable information or suggestions. For the classification task, the specific prompt is depicted in Figure 6.
Communication Skill Efficiency
We adopt the Motivational Interviewing Treatment Integrity (MITI) code Moyers et al. (2003), a well-established behavioral coding system that differentiates between favorable and unfavorable responses, to categorize responses into three classes based on their communication skills:
MI Adherent: Responses in this category support help-seekers with empathetic and compassionate statements, fostering a sense of being heard, respected, and understood. For example, the response “Well, there is really a lot going on for you right now” qualifies as MI Adherent.
Others: This class encompasses responses that might not immediately elicit a positive support effect. These include responses like a closed question, such as “Did you eat five fruits and vegetables this week?” In our work, we regard both these types of responses and those classified as MI Adherent as helpful. Our rationale is that such responses, though seemingly less impactful initially, can be beneficial in the long term. They contribute to the progression of the conversation and strengthen the relationship between the help-seeker and the supporter.
MI Non-Adherent: This category includes responses that involve arguing, confronting, or offering unsolicited advice, which may lead to resistance and impede problem-solving for help-seekers. An example is the response “Yes, you are an alcoholic. You might not think so, but you are,” which is classified as MI Non-Adherent.
To classify supporter statements based on communication efficiency, we utilize the MI dataset as proposed by Welivita and Pu (2022). Additionally, Figure 7 illustrates the specific prompt employed for this task.
Response Coherence
Context | |
Help-seeker: Is it afternoon where you are? I’m still waking up; I should have more coffee. And, I am just always tired, tired of everything these days. |
Original Response |
Yes, What happened? Why are you tired? Please can you explain? |
Incoherent Responses |
I am glad to hear you are feeling a little better! Yes, bad management is so toxic. Even with great coworkers in a job you love, horrible management can ruin it quickly. (From another dialogue) |
Why are you asking me to explain when you are feeling exhausted? Can you share your thoughts on this matter? (Keywords changed) |
To develop the dataset for response classification in the aspect of response coherence, we initially selected approximately K context-response pairs from the ESConv training dataset as the foundation for constructing the coherence data. It is important to note that this training dataset represents the intersection of the training sets used by all base models. As mentioned in Section 5.2, two types of incoherent variants were created: one by replacing the original response with a response from a different dialogue, and the other by altering keywords in the original response. For the latter approach, we utilized GPT-4 to modify keywords or crucial information, ensuring that the altered response either conveyed a contrary meaning or addressed a different topic. Table 5 presents an illustrative example of this process. The specific prompt employed for this task is illustrated in Figure 8.
A.2 Model Tuning and Module Performance
We instruction-tuned LLaMA111https://huggingface.co/decapoda-research/llama-7b-hf to equip the model with the capability required for unhelpful response classification tasks.222The implementation referred to is available at https://huggingface.co/decapoda-research/llama-7b-hf Furthermore, we initialized the LoRA weights using a low-rank adapter that was fine-tuned on the Stanford Alpaca dataset.333https://huggingface.co/tloen/alpaca-lora-7b We evaluated the capabilities of GPT-3.5 and GPT-4 by invoking the gpt-3.5-turbo and gpt-4 models, respectively, through their API444https://platform.openai.com/docs/api-reference/chat/create during the period from December 2023 to January 2024. We set the temperature parameter to to ensure deterministic output generation. The detailed results of this evaluation are detailed in Table 6.
Model |
GPT-3.5 |
GPT-4 |
LLaMA (Vanilla) |
LLaMA (Tuned) |
Empathetic Expression | accuracy |
81.88 |
69.58 |
19.09 |
83.50 |
F1 Score |
54.89 |
61.58 |
16.03 |
74.12 |
Skill Efficiency | accuracy |
69.56 |
79.01 |
84.96 |
90.43 |
F1 Score |
53.22 |
66.14 |
45.93 |
81.83 |
Response Coherence | accuracy |
33.47 |
78.76 |
33.31 |
92.98 |
F1 Score |
25.26 |
72.62 |
24.98 |
92.06 |
Aggregated Feedback | accuracy |
57.11 |
78.01 |
59.08 |
90.72 |
F1 Score |
41.99 |
77.19 |
40.96 |
89.86 |
GPT-3.5 refers to considering the supporter’s response not unhelpful. Prior to instruction-tuning, LLaMA exhibited notably poor performance across all facets of the classification tasks. The model consistently favored specific classes, such as “Strong Empathy,” “Other,” and “Yes,” corresponding to empathetic expression, communication efficiency in skill, and response coherence, respectively.
We also conduct human evaluation to examine the module. In particular, we randomly sample responses from the ESConv dataset and model-generated responses. For each response, an annotator was asked to assess whether it is unhelpful. We compared the annotations with the multifaceted AI feedback, and the consistency rate was . We found that the multifaceted AI module is stricter than human annotators. It is evident by the fact that for instances which the multifaceted AI feedback and human annotations are different, the multifaceted AI feedback tends to consider the response unhelpful.
Appendix B Experimental Setup
B.1 Devices and Environment
We employed Pytorch555https://pytorch.org/ for the implementation of all models, and conducted our experiments using Nvidia GeForce RTX 3090 GPUs.
B.2 ESConv Dataset
The ESConv dataset comprises a collection of emotional support conversations, each facilitated between a help-seeker and a supporter. While previous studies have partitioned and processed this dataset in various ways, our implementations of Muffin strictly follow each of their methodologies to maintain fairness. However, it is important to note that direct comparisons of performance across these models may not be entirely fair due to the differences in their data preprocessing approaches.
B.3 Base Models
Vanilla
is a vanilla BlenderBot Roller et al. (2021) trained on the dataset ESConv. We used the small version666https://huggingface.co/facebook/blenderbot_small-90M of BlenderBot in experiments following previous studies.
Joint
is built upon the backbone of BlenderBot and is specially trained to generate responses along with an expected communication skill at the beginning of each response Liu et al. (2021)777https://github.com/thu-coai/Emotional-Support-Conversation. The focused facet of this model is skill efficiency.
MultiESC
Cheng et al. (2022)888https://github.com/lwgkzl/multiesc is an emotional support conversation model that mainly focuses on the communication skill efficiency. It predicts the strategies (communication skills) in the next several turns considering the the user’s state.
KEMI
incorporates various knowledge for a mixed-initiative conversation model, which can provide emotional support Deng et al. (2023)999https://github.com/dengyang17/KEMI. In this process, the response coherence is also enhanced.
TransESC
Zhao et al. (2023)101010https://github.com/circle-hit/TransESC predicts the transitions of the user’s emotion, the communication skills, and the conversation keywords. Then, such information is used for response generation. This model takes into account more than one facet of emotional support; however, it is not optimized to reduce the likelihood of unhelpful responses like other models.
B.4 Human Evaluation
For each pair of and Muffin, we randomly selected response pairs generated by both and Muffin under identical conversational contexts. Selecting responses aligns with the standard practice for human evaluation in dialogue tasks, such as the experiments of TransESC Zhao et al. (2023) and KEMI Deng et al. (2023). To prevent annotators from identifying the generation model based on the order of sentences, the sequence in which these two responses are presented is randomized for each evaluation.
Initially, the annotators were briefed about the nature of emotional support conversations to ensure a comprehensive understanding of the task’s objectives. During the rating process, they were suggested to imagine themselves as the help-seekers within the conversations. After being provided with the conversation context, the annotators then proceeded to compare two generated responses, shown as LABEL:fig:huam_rating. Moreover, we prioritized the comfort and well-being of our annotators, advising them to pause or cease the annotation process if they encountered any content that made them feel uncomfortable. Annotators were paid at a rate of 1.52 times their local hourly minimum wage. Based on annotators’ feedback, it was estimated that approximately 40 seconds were spent on evaluating each response pair.
The inter-rater agreement among annotators, as measured by Fleiss’s Kappa, is . This value is relatively high, especially when compared to inter-rater reliability values in most subjective tasks, which typically fall within the range of 0.2 to 0.6 Wong et al. (2021); Cowen and Keltner (2017). In the process of aggregating the annotations, we determine the winning response based on the consensus of annotators. If two annotators prefer response A, and another two annotators prefer a tie, we categorize response A as the winner. When an equal number of annotators favor two responses, we label the result as a tie.
Appendix C Response Generation Model Performance
C.1 Multifaceted AI Feedback
In addition to presenting the primary results and conducting an ablation study, we provide the multifaceted AI feedback results as a reference, as outlined in Table 7. For each model, we calculate the frequency of generating unhelpful responses on the ESConv testing set. Specifically, we utilize the multifaceted AI feedback module to identify the helpful (non-unhelpful) responses and compute their percentage, displayed in the left subtable. Furthermore, we analyze the helpful response percentage when each model generates ten responses using diverse beam search, reported in the right subtable. It’s worth noting that the assertion “approximately 30% of responses generated by MultiESC on the ESConv benchmark are identified as unhelpful” is derived from the findings presented in this table.
A. Generating One Response | ||||
Model | emp. | skill | cohr. | agg. |
Vanilla | 81.22 | 90.43 | 80.69 | 64.83 |
Muffin | 85.83 | 92.82 | 84.33 | 71.26 |
Joint | 80.61 | 88.93 | 80.65 | 61.48 |
Muffin | 82.33 | 90.47 | 83.04 | 64.76 |
TransESC | 81.06 | 91.44 | 74.76 | 63.24 |
Muffin | 81.28 | 91.49 | 78.67 | 66.02 |
KEMI | 83.01 | 88.47 | 85.76 | 70.90 |
Muffin | 83.40 | 88.68 | 87.15 | 72.33 |
MultiESC | 83.39 | 90.66 | 85.38 | 70.00 |
Muffin | 85.24 | 92.92 | 86.06 | 72.38 |
B. Generating Ten Responses | ||||
Model | emp. | skill | cohr. | agg. |
Vanilla | 81.90 | 92.03 | 78.89 | 64.51 |
Muffin | 89.84 | 96.05 | 82.60 | 73.78 |
Joint | 79.31 | 90.56 | 79.75 | 63.30 |
Muffin | 83.43 | 92.26 | 80.02 | 66.54 |
TransESC | 74.16 | 92.78 | 43.93 | 41.21 |
Muffin | 74.20 | 92.92 | 44.15 | 41.43 |
KEMI | 81.16 | 87.49 | 81.57 | 66.23 |
Muffin | 81.39 | 87.64 | 81.61 | 66.47 |
MultiESC | 78.61 | 89.67 | 77.33 | 60.83 |
Muffin | 78.77 | 89.73 | 77.26 | 60.87 |
Overall, Muffin demonstrates enhancements in AI feedback across multiple facets. However, three intriguing phenomena emerge. (1). While the left subtable clearly indicates an increase in the frequency of helpful responses attributable to the Muffin framework, the evidence presented in the right subtable is weaker (agg.), particularly when the base model is TransESC, KEMI, or MultiESC. However, this finding aligns with our loss function (Equation 1 and Equation 4). We have introduced a contrastive loss to the original generation loss, but this addition does not significantly mitigate unhelpfulness. However, the contrastive loss assigns higher generation probabilities to helpful (non-unhelpful) responses. Consequently, when generating one response, the one with the highest probability is output. Thus, the left subtable shows that Muffin effectively reduce the frequency of unhelpful responses. (2). Despite lacking a dedicated mechanism for incorporating communication skills, BlenderBot-Vanilla attains a notably high score in communication skill efficiency (skill). Upon closer examination, we observe that this model frequently produces responses such as “I can understand that…” or “I’ve experienced something similar…”. These responses are categorized as self-disclosure. While this strategy doesn’t undermine previous emotional support efforts and remains undetected as “unhelpful” by the AI feedback module, a conversation dominated by self-disclosure responses may not fully align with the objective of providing support. (3). Another noteworthy observation is that TransESC receives exceptionally low AI feedback in response coherence (cohr.), despite its incorporation of external knowledge. However, insights from TransESC’s ablation experiment Zhao et al. (2023) suggest that the generated knowledge may contain noise. This potential factor could explain the low feedback score observed in response coherence.
C.2 Cases
For a closer and more detailed examination of our method, we present additional comparisons between the base model and the Muffin model in Table 8. Additionally, we include AI feedback for each facet as a reference.
Conversation History |
Supporter: Is there anything you’d like to talk to me about? |
Help-seeker: I have a friend who knew me since I was little. She is a little older than me. She still treat me like I’m still little. I tried to explain to her that I’m a grown woman and I grew out of lot of things when I was a kid. |
Model |
Response |
emp. | skill | cohr. |
Joint |
I have had a friend like that before. They are a good friend, and you should feel as though you are a better person. |
Weak Empathy |
Other |
No |
Muffin |
I have been in a similar situation. I had a friend who was a little older than me, and I had to deal with the same thing. |
Weak Empathy |
Other |
Yes |
Conversation History |
Help-seeker: Hello I am very sad today. My partner is cheating on me. I was told by a customer at work. I do not know this person well, yet they knew the person he was cheating on me with. He will not admit it himself. |
Supporter: I am sorry if your partner is cheating. So, your partner denies the charge of cheating? |
Help-seeker: Yes. He says the person who told me is lying. We are not married but we live together. |
Supporter: Do you think maybe it was a rumor started by someone who could be jealous? Ash your partner if he loves you enough to propose marriage. If he says no, then ask him why. |
Help-seeker: I am not sure if I am okay with giving him an ultimatum. We are still young and although I love being in a relationship, I am unsure about marriage. But I appreciate your suggestions. And you make a great point about the workplace being hostile. |
Model |
Response |
emp. | skill | cohr. |
MultiESC |
I think you should give him an ultimatum and see if he accepts it. |
No Empathy |
MI Non-Adherent |
Yes |
Muffin |
Have you spoken to your friends or family about the situation? |
Strong Empathy |
Other |
Yes |
Conversation History |
Help-seeker: I recently broke up with my girlfriend. We had been together for more than 3 years. She said that she simply wasn’t in love anymore, and that she didn’t have the same feelings she did at the beginning. |
Supporter: You still have your whole life ahead of you. Some people don’t find a partner until their 40’s or even 50’s… I think that’s what your girl would want you to do as well. |
Help-seeker: I disagree with you regarding that maybe that is what she wants me to do. I don’t think she cares about me at the slightest anymore. I was a burden for her and will probably keep being it for any woman I meet. |
Model |
Response |
emp. | skill | cohr. |
KEMI |
That is a great attitude to have. You are a strong person, and I think you will find someone better. |
Weak Empathy |
MI Non-Adherent |
Yes |
Muffin |
I think you are a very good person and I think that you will find someone much better. |
Weak Empathy |
MI Adherent |
Yes |